# Direct Imaging of Hydrogen-Driven Dislocation and Strain Field Evolution in a Stainless Steel Grain

### v2

@@ 2.3 Dislocation dynamics
A phase vortex in a BCDI reconstruction indicates the spatial position of a dislocation core [43], enabling the mapping of the 3D dislocations (Methods). Figure 3 shows the dislocation evolutions during hydrogen charging. Initially, the grain shows two small dislocations close to the surface, and one large bow-shaped dislocation that extends across the grain, hereafter referred to as the “large dislocation.” The large dislocation remains unchanged before hydrogen charging and up to 1.9 h after the start of charging. To determine its Burgers vector, we assume that it is a shear loop and lies in the $\left(\bar{1}11\right)$ plane, as expected for FCC metals (we confirm the angle between the dislocation glide plane normal and $\mathbf{Q}_{111}$ is 68.9°, less than 2° from the angle between theoretical $\langle 111\rangle$ directions and within error). Thus, we determine the orientation matrix, $\mathbf{UB}$ (Supporting Text). Glide dislocations in the $\left(\bar{1}11\right)$ plane can have three different Burgers vectors $\mathbf{b}=\frac{a_{0}}{2}[110]$, $\frac{a_{0}}{2}[101]$ or $\frac{a_{0}}{2}[0\bar{1}1]$. However, $\mathbf{Q}_{111}\cdot\mathbf{b}\neq 0$ only for $\frac{a_{0}}{2}[110]$ and $\frac{a_{0}}{2}[101]$, meaning that only dislocations with these Burgers vectors will be visible in a $\mathbf{Q}_{111}$ crystal reflection [48]. By comparing the measured strain field of the large dislocation to a 3D elastic model and dislocation dynamics simulations, we determine that the large dislocation has Burgers vector $\mathbf{b}=\frac{a_{0}}{2}[110]$ (Supporting Text, Figs. LABEL:supp-fig:dislocation_model_slices-LABEL:supp-fig:dislocation_dynamics_models, and Table LABEL:supp-table:error).
@@
Figure 3: Evolution of the dislocations before and during hydrogen charging. a,b,c), Columns represent a different orthogonal view of the grain and the dislocations. The top row shows the initial morphology of the grain rendered as a grey isosurface. The subsequent rows have translucent renderings of the same grain morphology (largely unchanged, see Figure 2) along with the dislocations at different states, with glide and climb events indicated. The dislocation dynamics simulation of glide is indicated in green. Dislocations are colored according to the time of observation. The black arrow indicates the Burgers vector. See Supporting Video S2 for the dislocations at each measurement time.
@@
After 1.9 h of hydrogen charging, the large dislocation begins to glide (Figure 3). This is remarkable, as Figure 2 shows little change in surface strain, and hence externally-applied stress that could drive plasticity. It suggests that hydrogen either reduces the Peierls barrier for dislocation glide, or unpins the dislocation from obstacles such as alloying elements [54]. Either way, the introduction of hydrogen enables the dislocation to relax, reducing its line length and thus strain energy (Supporting Text and Figure LABEL:supp-fig:dislocation_elastic_energy) [53]. A discrete dislocation dynamics simulation (Methods) is set up with the large dislocation in the position observed before hydrogen charging. The ends of the dislocation are pinned at the grain boundary, consistent with our experimental observation, and the system was then allowed to relax under self-stress. The subsequent dislocation evolution, shown in Figure 3, agrees very well with that observed experimentally. This confirms that the dislocation glide observed in experiments is relaxation-driven, suggesting hydrogen-enhanced dislocation mobility and/or depinning.
@@
Following the glide motion, after 3.2 h of hydrogen charging, one end of the large dislocation becomes unpinned and climbs out of the glide plane. Subsequently, from 3.9 h onwards, the dislocation remains largely unchanged upon further charging. The climb force on a mixed dislocation is the sum of an elastic interaction force and an osmotic force [54]. Figure 2c shows that the average strain on the grain surface changes little, indicating that the contribution of external stresses to dislocation climb will be small. Furthermore, the unpinned dislocation segment is $\sim 200$ nm from the nearest observable small dislocation, meaning that any dislocation-dislocation forces will be minor [16]. Thus, the climb force on the large dislocation will be dominated by the osmotic force, $F_{\mathrm{os}}$. $F_{\mathrm{os}}$, per unit length of a dislocation line is [54],
@@
Our observations suggest that once there is a sufficient hydrogen concentration within the grain, there will be enough vacancies and sufficient reduction in vacancy and metal migration energy such that the osmotic force can facilitate climb. The equilibrium concentration of vacancies along the length of the large dislocation depends on the spatial variation of hydrogen concentration. Based on the magnitudes of $e_{111}$ $\left(\sim 10^{-3}\right)$ compared to $\varepsilon_{111}$ $\left(\sim 10^{-4}\right)$, increasing hydrogen concentration dominates homogeneous lattice swelling. Therefore, any variations in hydrogen concentration along the dislocation, and by extension, vacancy concentration, are small by comparison. This suggests that differences in $b_{e}$, i.e. the degree of edge character along the dislocation line, will determine which parts of the dislocation climb. Figure 4 shows a magnified view of the large dislocation before and after the climb at 3.9 h. Each segment of the dislocation is colored by edge character. The edge character is low at the end of the dislocation that does not climb, making this segment less likely to climb due to a lower $F_{\mathrm{os}}$. Conversely, the segment of the dislocation that climbs also has the greatest edge character. The climb direction (magenta arrow in Figure 4) is nearly perpendicular (89.5°) to the Burgers vector, as expected. After climb, the climbing segment lies on the $\left(11\bar{1}\right)$ plane. However, since $\mathbf{b}=\frac{a_{0}}{2}[110]$ is not one of the Burgers vectors in this plane, this segment is sessile, explaining the lack of further evolution beyond 3.9 h.
Dislocation climb is surprising since dislocation cross-slip is more energetically favored. Dislocation cross-slip requires dislocation screw character. Here, the Burgers vector is $\mathbf{b}=\frac{a_{0}}{2}[110]$, which is not parallel to the dislocation line segment that unpins, thus the segment cannot cross-slip from the $\left(\bar{1}11\right)$ plane to the $\left(11\bar{1}\right)$ plane (Figure 4). Unfortunately, we do not know what the dislocation looks like between 3.2 h and 3.9 h, so we cannot rule out cross-slip based on the absence of screw character alone. However, cross-slip also requires high stacking fault energy, which is known to be reduced by hydrogen in austenitic SS [58]. It also requires applied stress, which is minimal based on Figure 2. Thus, we propose that vacancy-assisted climb is most plausible in the hydrogen environment.
@@
Figure 4: Dislocation unpinning and climb. A translucent isosurface of the grain showing only the large dislocation before (1 - at 3.2 h) and after the climb event (2 - at 3.9 h). The dislocation is colored between two nodes based on the $b_{e}$ (Equation 2), where $b_{e}/|\mathbf{b}|=1$ represents a pure edge dislocation and $b_{e}/|\mathbf{b}|=0$ represents a pure screw dislocation. The black arrow indicates the Burgers vector direction. The magenta arrow, which is nearly perpendicular to the Burgers vector direction, indicates the direction of climb. a), Orthogonal views based on sample coordinates. b), Crystallographic views based on the dislocation loop plane, with the dislocation initially lying on the $\left(\bar{1}11\right)$ plane, and later one end unpinning and climbing onto the $\left(11\bar{1}\right)$ plane.
@@ 2.4 Dislocation strain field evolution
Based on observations of dislocation pileup evolution upon introduction of hydrogen, it has been previously inferred that hydrogen reduces the interaction between dislocations by changing their stress fields [16, 41]. To quantify the changes in the strain field surrounding the large dislocation, we compare slices though the 3D reconstructed volume where $\mathbf{b}$ and $[\bar{1}11]$ (the normal to the initial glide plane of the large dislocation) are in plane and $\xi$ is normal to the slice (i.e. where the large dislocation has predominantly edge character). These slices are shown in Figure 5a and b. A theoretical model, devoid of hydrogen, of an edge dislocation strain field projected along $[111]$ is shown in Figure 5c (Methods), and compared to the experimental strain field by considering circular line profiles drawn at a 30 nm radius from the dislocation core (Figure 5d). Agreement between the theoretical model and the experiment is initially good, but worsens as the hydrogen concentration increases.
@@
Figure 5: Evolution of the internal strain field surrounding the large dislocation. Three time points are presented as rows in a,b,c,d). a), Translucent morphologies of the grain with dislocations. Slices through $\varepsilon_{111}$ are in-plane to the Burgers vector and the normal to the $\left(\bar{1}11\right)$ plane. The average lattice parameter is listed for each time. b), A $\varepsilon_{111}$ slice capturing a section of the large dislocation with pure edge character. See Supporting Video S3 for $\varepsilon_{111}$ slices through the entire grain. c), Theoretical elastic model, $\varepsilon_{111,\mathrm{model}}$, of the large dislocation devoid of hydrogen (Methods). d), Circular line profiles, drawn at a 30 nm radius from each dislocation core, compared to the model. The shaded region is the standard deviation of the experimental values in the neighboring 26 pixels. Line profiles start at the Burgers vector (horizontal) and run anticlockwise. e), Circular line profiles drawn for each time point of the experiment, plotted as a surface indicating $\varepsilon_{111}$ values. f), Top-down view of e). g), Maximum and minimum values of the line profiles, averaged over a range of $\pi/4$. The shaded region corresponds to one standard deviation of each value. The model values are offset to overlap with the experimental values at negative times.
@@
The evolution of strain along the circular line profiles during charging is shown in Figure 5e and f. Figure 5g shows how the maximum and minimum values, averaged over ±$\pm\pi/4$ around each maximum/minimum, evolve during charging. We observe a reduction of the elastic strain surrounding the dislocation by up to 35% between the start and end of hydrogen charging. This relative reduction agrees well with the computed reduction of edge dislocation-carbon atom interaction energy in $\alpha$-Fe upon the introduction of hydrogen [59].
@@
At the nanoscale, we attribute this elastic strain reduction to the formation of a hydrogen atmosphere in the vicinity of the dislocation. Dislocations are relatively large and strong hydrogen trapping sites ($\sim 27\ \mathrm{kJ\ mol^{-1}}$) [22], so the hydrogen atoms surrounding the dislocations are relatively immobile at room temperature [24]. Therefore, the evolution of this localized accumulation can be probed using the heterogeneous strain. Figure 5g also shows that the maximum $\varepsilon_{111}$ deviates more from the elastic hydrogen-free model than the minimum $\varepsilon_{111}$ This suggests that hydrogen prefers to reside in the tensile region of the dislocation [16]. Alternatively, the reduction in tensile strain could also be due to the hydrogen-facilitated diffusion of solutes, resulting in solute segregation [53], but this is unlikely since we are at room temperature.
@@ 2.5 Proposed mechanisms
Our results capture the evolution of a single dislocation undergoing glide relaxation, followed by climb. We propose that this sequence is initiated by the introduction of hydrogen into the lattice, which greatly enhances dislocation mobility via several complementary mechanisms. A hydrogen-induced reduction of lattice friction, or Peierls stress, has been predicted by DFT studies on pure metals using the Peierls-Nabarro model [60]. However, in 316 SS, this reduction is small compared to the pinning stress from alloying elements, given that 316 SS’s yield stress is more than twice that of pure iron. In addition, hydrogen accumulation has been predicted to lower the dislocation core energy [61] and reduce the dislocation strain field [59], as we have directly measured (Figure 5). This reduces the elastic interactions of the dislocation with the surrounding environment, such as other dislocations and point defects, allowing easier dislocation glide. Furthermore, our observation that hydrogen greatly enhances dislocation climb mobility (Figure 4) suggests that also during glide, dislocations will be able to “climb around” obstacles that would otherwise lead to dislocation pinning. Together, these observations suggest that, in the presence of hydrogen, the currently accepted view of glide-dominated dislocation motion [16, 14, 41] may no longer hold.

### v4

@@ 2.3 Dislocation dynamics
A phase vortex in a BCDI reconstruction indicates the spatial position of a dislocation core [46], enabling the mapping of 3D dislocations (Experimental Section). Figure 3 shows the dislocation evolution during hydrogen charging. Initially, the grain shows two small dislocations close to the surface, and one large bow-shaped dislocation that extends across the grain, hereafter referred to as the “large dislocation.” The large dislocation remains unchanged before hydrogen charging and up to 1.9 h after the start of charging. To determine its Burgers vector, we assume that it is a shear loop and lies in the $\left(\bar{1}11\right)$ plane, as expected for FCC metals (we confirm the angle between the dislocation glide plane normal and $\mathbf{Q}_{111}$ is 68.9°, less than 2° from the angle between theoretical $\langle 111\rangle$ directions and within error). Thus, we determine the orientation matrix, $\mathbf{UB}$ (Supporting Information). Glide dislocations in the $\left(\bar{1}11\right)$ plane can have three different Burgers vectors $\mathbf{b}=\frac{a_{0}}{2}[110]$, $\frac{a_{0}}{2}[101]$ or $\frac{a_{0}}{2}[0\bar{1}1]$. However, $\mathbf{Q}_{111}\cdot\mathbf{b}\neq 0$ only for $\frac{a_{0}}{2}[110]$ and $\frac{a_{0}}{2}[101]$, meaning that only dislocations with these Burgers vectors will be visible in a $\mathbf{Q}_{111}$ crystal reflection [51]. By comparing the measured strain field of the large dislocation to a 3D elastic model and discrete dislocation dynamics (DDD) simulations, we determine that the large dislocation has Burgers vector $\mathbf{b}=\frac{a_{0}}{2}[110]$ (Supporting Information, Figures LABEL:supp-fig:dislocation_model_slices and LABEL:supp-fig:dislocation_dynamics_models, and Table LABEL:supp-table:error, Supporting Information).
@@
Figure 3: Evolution of the dislocations before and during hydrogen charging. a,b,c) Columns represent a different orthogonal view of the grain and the dislocations. The top row shows the initial morphology of the grain rendered as a grey isosurface. The subsequent rows have translucent renderings of the same grain morphology (largely unchanged, see Figure 2) along with the dislocations at different states, with glide and climb events indicated. The dislocation dynamics simulation of glide is indicated in green. Dislocations are colored according to the time of observation. The black arrow indicates the Burgers vector. See Video S2 (Supporting Information) for the dislocations at each measurement time.
@@
After 1.9 h of hydrogen charging, the large dislocation begins to glide (Figure 3). This is remarkable, as Figure 2 shows little change in surface strain, and hence externally applied stress that could drive plasticity. It suggests that hydrogen either reduces the Peierls barrier for dislocation glide, or unpins the dislocation from obstacles such as alloying elements [57]. Either way, the introduction of hydrogen enables the dislocation to relax, reducing its line length and thus strain energy (Supporting Information and Figure LABEL:supp-fig:dislocation_elastic_energy, Supporting Information) [56]. A DDD simulation (Experimental Section) is setup with the large dislocation in the position observed before hydrogen charging. The ends of the dislocation are pinned at the grain boundary, consistent with our experimental observation, and the system was then allowed to relax under self-stress. The subsequent dislocation evolution, shown in Figure 3, agrees well with that observed experimentally. This confirms that the dislocation glide observed in experiments is relaxation-driven, suggesting hydrogen-enhanced dislocation mobility and/or depinning.
@@
Following the glide motion, after 3.2 h of hydrogen charging, one end of the large dislocation becomes unpinned and climbs out of the glide plane. Subsequently, from 3.9 h onward, the dislocation remains largely unchanged upon further charging. The climb force on a mixed dislocation is the sum of an elastic interaction force and an osmotic force [57]. Figure 2c shows that the average strain on the grain surface changes little, indicating that the contribution of external stresses to dislocation climb will be small. Furthermore, the unpinned dislocation segment is $\approx 200$ nm from the nearest observable small dislocation, meaning that any dislocation–dislocation forces will be minor [19]. Thus, the climb force on the large dislocation will be dominated by the osmotic force, $F_{\mathrm{os}}$. $F_{\mathrm{os}}$, per unit length of a dislocation line is [57],
@@
Our observations suggest that once there is a sufficient hydrogen concentration within the grain, there will be enough vacancies and sufficient reduction in vacancy and metal migration energy such that the osmotic force can facilitate climb. The equilibrium concentration of vacancies along the length of the large dislocation depends on the spatial variation of hydrogen concentration. Based on the magnitudes of $e_{111}$ $\left(\approx 10^{-3}\right)$ compared to $\varepsilon_{111}$ $\left(\approx 10^{-4}\right)$, increasing hydrogen concentration dominates homogeneous lattice swelling. Therefore, any variations in hydrogen concentration along the dislocation, and by extension, vacancy concentration, are small by comparison. This suggests that differences in $b_{e}$, i.e. the degree of edge character along the dislocation line, will determine which parts of the dislocation climb. Figure 4 shows a magnified view of the large dislocation before and after the climb at 3.9 h. Each segment of the dislocation is colored by edge character. The edge character is low at the end of the dislocation that does not climb, making this segment less likely to climb due to a lower $F_{\mathrm{os}}$. Conversely, the segment of the dislocation that climbs also has the greatest edge character. The climb direction (magenta arrow in Figure 4) is nearly perpendicular (89.4°) to the Burgers vector, as expected.
After climb, the climbing segment lies on the $\left(11\bar{1}\right)$ plane. However, since $\mathbf{b}=\frac{a_{0}}{2}[110]$ is not one of the Burgers vectors in this plane, this segment is sessile, explaining the lack of further evolution beyond 3.9 h. This sudden dislocation pinning may result from a sufficiently high hydrogen concentration accumulating at the dislocation core [61], as recently observed in ferritic steel using ECCI [62]. Alternatively, high concentrations of hydrogenated vacancies could lead to dislocation immobilization, as shown in other in situ TEM experiments on initially mobile dislocations in aluminum [63].
[…]
@@
Figure 4: Dislocation unpinning and climb. A translucent isosurface of the grain showing only the large dislocation before (1, at 3.2 h) and after the climb event (2, at 3.9 h). The dislocation is colored between two nodes based on the $b_{e}$ (Equation 2), where $b_{e}/|\mathbf{b}|=1$ represents a pure edge dislocation and $b_{e}/|\mathbf{b}|=0$ represents a pure screw dislocation. The black arrow indicates the Burgers vector direction. The magenta arrow, which is nearly perpendicular to the Burgers vector direction, indicates the direction of climb. a) Orthogonal views based on sample coordinates. b) Crystallographic views based on the dislocation loop plane, with the dislocation initially lying on the $\left(\bar{1}11\right)$ plane, and later one end unpins and climbs onto the $\left(11\bar{1}\right)$ plane.
@@ 2.4 Dislocation strain field evolution
Based on observations of dislocation pileup evolution upon introduction of hydrogen, it has been previously inferred that hydrogen reduces the interaction between dislocations by changing their stress fields [19, 44]. To quantify the changes in the strain field surrounding the large dislocation, we compare slices though the 3D reconstructed volume where $\mathbf{b}$ and $\bar{1}11]$ (the normal to the initial glide plane of the large dislocation) are in plane and $\xi$ is normal to the slice (i.e. where the large dislocation has predominantly edge character). These slices are shown in Figure 5a, b. A theoretical model, devoid of hydrogen, of the edge dislocation strain field projected along $[111]$ is shown in Figure 5c (Experimental Section), and compared to the experimental strain field by considering circular line profiles drawn at a 30 nm radius from the dislocation core (Figure 5d). Agreement between the theoretical model and the experiment is initially good, but worsens as the hydrogen concentration increases.
@@
Figure 5: Evolution of the internal strain field surrounding the large dislocation. Three time points are presented as rows in (a–d). a) Translucent morphologies of the grain with dislocations. Slices through $\varepsilon_{111}$ are in-plane to the Burgers vector and the normal to the $\left(\bar{1}11\right)$ plane. The average lattice parameter is listed for each time. b) A $\varepsilon_{111}$ slice capturing a section of the large dislocation with pure edge character. See Video S3 (Supporting Information) for $\varepsilon_{111}$ slices through the entire grain. c) Theoretical elastic model, $\varepsilon_{111,\ \mathrm{model}}$, of the large dislocation devoid of hydrogen (Experimental Section). d) Circular line profiles, drawn at a 30 nm radius from each dislocation core, compared to the model. The shaded region is the standard deviation of the experimental values in the neighboring 26 pixels. Line profiles start at the Burgers vector (horizontal) and run anticlockwise. e) Circular line profiles drawn for each time point of the experiment, plotted as a surface indicating $\varepsilon_{111}$ values. f) Top-down view of (e). g) Maximum and minimum values of the line profiles, averaged over a range of $\pi/4$. The shaded region corresponds to one standard deviation of each value. To better highlight the differences between the model and experimental data at positive charging times, the model’s maximum and minimum values were slightly offset to align with the experimental data at negative times.
@@
The evolution of strain along the circular line profiles during charging is shown in Figure 5e,f. Figure 5g shows how the maximum and minimum values, averaged over $\pm\pi/4$ around each maximum/minimum, evolve during charging. We observe a reduction of the elastic strain surrounding the dislocation by up to 35% between the start and end of hydrogen charging. This relative reduction agrees well with the computed reduction of edge dislocation-carbon atom interaction energy in $\alpha$-Fe upon the introduction of hydrogen [67].
@@
At the nanoscale, we attribute this elastic strain reduction to the formation of a hydrogen atmosphere in the vicinity of the dislocation. Dislocations are relatively large and strong hydrogen trapping sites ($\approx 27\ \mathrm{kJ\ mol^{-1}}$) [25], so the hydrogen atoms surrounding the dislocations are relatively immobile at room temperature [27]. Therefore, the evolution of this localized accumulation can be probed using the heterogeneous strain. Figure 5g also shows that the maximum $\varepsilon_{111}$ deviates more from the elastic hydrogen-free model than the minimum $\varepsilon_{111}$. This suggests that hydrogen prefers to reside in the tensile region of the dislocation [19]. Alternatively, the reduction in tensile strain could also be due to the hydrogen-facilitated diffusion of solutes, resulting in solute segregation [56], but this is unlikely since we are at room temperature.
@@ 2.5 Proposed mechanisms
We now consider whether dislocation mobility can be introduced by internal stresses induced by hydrogen concentration gradients caused by cathodic hydrogen charging [68]. As hydrogen diffuses into the SS disk from the top, lattice swelling ensues [54, 29], creating a hydrostatic internal stress. There is an in-plane boundary constraint from the uncharged bulk that generates an in-plane stress on the grain, such that the in-plane strain remains zero, as required for compatibility [69]. We have modeled the experimental conditions and predicted the hydrogen concentration as a function of time and depth below the surface of the disk (Supporting Information and Figure LABEL:supp-fig:H_concentration_profilea, Supporting Information). The resulting strain distributions can be compared to the experimental data. In Fig. LABEL:supp-fig:Strain_comparison, Supporting Information, we observe that there is an additional background strain field in the measured $\varepsilon_{111}+e_{111}$ data that is not directly captured by the hydrogen-induced strain gradient combined with the large dislocation strain field model, $\varepsilon_{111,\ \mathrm{model}}+\varepsilon_{111,\ \mathrm{b.c.}}$. We do note that this background remains largely unchanged throughout the charging, suggesting that stresses due to neighboring grains remain largely unchanged during hydrogen uptake.
The hydrogen uptake creates a resolved shear stress that acts on the dislocation. The in-plane stress has been applied in the DDD model, which shows only a slight lift in part of the dislocation toward the unpinning direction (orange arrows in Figure LABEL:supp-fig:dislocation_dynamics_models_stress, Supporting Information). Computation of the resulting Peach–Koehler (PK) force on the large dislocation reveals a significant component in the out-of-plane direction of dislocation climb (purple arrows in Figure LABEL:supp-fig:Fig4_bc, Supporting Information). Interestingly, the PK force also has an in-plane component opposite the direction of glide, which explains why the DDD model does not exhibit gliding behavior in Figure LABEL:supp-fig:dislocation_dynamics_models_stress, Supporting Information. This allows us to conclude that, while hydrogen gradient-induced stresses are significant, they do not appear to be the primary factor controlling dislocation glide. Instead, the hydrogen stress shielding effect [44] reduces the large dislocation’s effective Burgers vector, lowering the PK force and creating an energetically favorable condition where the dislocation line straightens, as we observe experimentally.
Our results capture the evolution of a single dislocation undergoing glide relaxation, followed by climb. We propose that this sequence is initiated by the introduction of hydrogen into the lattice, which greatly enhances dislocation mobility via several complementary mechanisms. A hydrogen-induced reduction of lattice friction, or Peierls stress, has been predicted by DFT studies on pure metals using the Peierls–Nabarro model [70]. However, in 316 SS, this reduction is small compared to the pinning stress from alloying elements, given that 316 SS’s yield stress is more than twice that of pure iron. In addition, hydrogen accumulation has been predicted to lower the dislocation core energy [71] and reduce the dislocation strain field [67], as we have directly measured (Figure 5). This reduces the elastic interactions of the dislocation with the surrounding environment, such as other dislocations and point defects, allowing easier dislocation glide. Furthermore, our observation that hydrogen greatly enhances dislocation climb mobility (Figure 4) suggests that also during glide, dislocations will be able to “climb around” obstacles that would otherwise lead to dislocation pinning. Together, these observations suggest that, in the presence of hydrogen, the currently accepted view of glide-dominated dislocation motion [19, 16, 44] may no longer hold.